OpenStudy (anonymous):

The Base of a prism is a regular pentagon with one side measuring 15cm. find the volume and the lateral surface area if the altitude measures 32cm.

OpenStudy (anonymous):

LSA=Pe where as:LSAis the lateral surface area, P is the perimeter of a right sector and e is the lenght of a lateral edge V=Bh where as V is volume and B is the area of the abse and h is the height of the prism V=Bh=Re where as r is the area of a right section and e is the length of a lateral edge, since Bh=Re and h +esin teta tghe R=Bsinteta

OpenStudy (anonymous):

this are the formulas

OpenStudy (nurali):

Here, side, a = 15 cm Area = a^2*sqrt[25 +10sqrt(5)] = = 15^2 * sqrt[25 +22.361] = 1174.23 cm^2 >======< ANSWER height = 32 cm VOLUME = 1174.23 * 32 = 37575.494 cm^3 >=====< ANSWER

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

but how did you get the area?